OpenStudy (pphalke):

If both legs of a triangle measure six inches, how long is the hypotenuse? Round your answer to the nearest tenth of an inch

OpenStudy (pphalke):

@lucifer @myininaya

OpenStudy (pphalke):

72

OpenStudy (medicaldoctor):

Can you show how you used the Pythagoras Theorem?

OpenStudy (medicaldoctor):

a^2 + b^2 = c^2

OpenStudy (medicaldoctor):

(6)^2 + (6)^2 = ?

OpenStudy (pphalke):

that's 36 +36 it equals 72

OpenStudy (medicaldoctor):

So (6)^2 is equal to 36 therefore 36 + 36 = 72

OpenStudy (medicaldoctor):

Yes that is correct

OpenStudy (pphalke):

it says to round to the nearest tenth of an inch

OpenStudy (medicaldoctor):

Okay well the nearest inch is 72 because it isn't a decimal number.

OpenStudy (agent0smith):

You forgot that it's c^2 = 72

OpenStudy (medicaldoctor):

Right exactly what I said.

OpenStudy (medicaldoctor):

I said that a^2 + b^2 = c^2

OpenStudy (lucifer):

http://prntscr.com/biac6e

OpenStudy (agent0smith):

@MedicalDoctor I meant when you said it's 72 inches, you forgot that it's still c^2 = 72.